OpenStudy (anonymous):

a car traveling at 85km/h slows down at a constant acceleration of .5m/s^2. Calculate the distance it coasts before it stops?

OpenStudy (anonymous):

85km/h=23.61m/s Again, use vf^2=vo^2 +2as \[s=0^{2}-23.61^{2}/2(-0.5)\] so s=557.4m

OpenStudy (anonymous):

i did but it came out as twice the amount :(